OpenStudy (anonymous):

Solve the equation. - 1/3m – 7 = 5 A. –4 B. –15 C. –36 D. 6

OpenStudy (anonymous):

Add 7. Then get -1/3m=12. Divide by -1/3 and get m=-4.

jhonyy9 (jhonyy9):

-1/3m -7= 5 -1 ---m -7 = 5 3 -1/3m=12 -m=36 m=-36

OpenStudy (anonymous):

I was thinking it was C.

OpenStudy (anonymous):

I Got 4.

jhonyy9 (jhonyy9):

C is right sure

jhonyy9 (jhonyy9):

do you understand it ?

OpenStudy (anonymous):

I think so. As best I can anyway. =)

jhonyy9 (jhonyy9):

good luck bye